OpenStudy (anonymous):

C=4(23245)^-1/3(13.5) Solve for C

OpenStudy (anonymous):

C=4(23245)^-1/3(13.5) C=4(0.35)(13.5) C=.14(13.5) C=1.89, But I'm trying to see how the 0.35 come from, 23245